nikenike:

Teresa’s age is 5 years more than twice Ben’s age. If Ben is x years old, the expression gives the sum of Ben’s and Teresa’s ages. If Ben is 7 years old, the sum of Ben’s and Teresa’s ages is years.

sophia333:

if Ben is 7 years old your going to do 7×2+5 to get Teresa's age then you add her age to his age. so your going to multiply 7×2 and add the answer of 7×2 to 5.

kacchan:

If Ben is x years old, then twice his age is 2x, and 5 years more than this is 2x+5 If Ben is x years old, then twice his age is 2x, and 5 years more than this is 2x+5 So, the sum of their ages is x+(2x+5)=3x+5 If Ben is 7 years old, i.e. if x=7, the sum becomes 3 * 7+5=21+5=26